# Bound states for systems with quartic energy-momentum dispersion

### v1

@@ Abstract
Bound states and its energies for systems with the quartic energy-momentum dispersion $E(p)\sim p^{4}$ and polynomial potentials are studied using the Wentzel-Kramers-Brillouin (WKB) semiclassical approximation and the Wentzel complex method taking into account higher order WKB corrections. The obtained energies are compared with numerical values found by applying the variational approach utilizing the universal Gaussian basis. It is found that the wave functions of the ground and higher-energy states for systems with quartic dispersion have nodes in the classically forbidden region. Thus, the well-known oscillation theorem for the one-dimensional Schrödinger equation is not, in general, applicable for systems with quartic dispersion. Still it is observed that the oscillation theorem holds in the classically allowed region in all considered examples. The properties of bound state wave functions are compared with the solutions of the exactly solvable problem of a square well potential.
@@ I Introduction
Quantum mechanical systems in which interactions or potential energy dominates over the kinetic energy attract a significant interest because they often exhibit strongly correlated behavior and give rise to novel quantum phases of matter which are not well described by traditional perturbative or mean-field approaches [1, 2]. These systems are central to many open questions in modern physics including entanglement, the simulation of strongly interacting gauge theories in condensed matter analogs, and the realization of quantum criticality and emergent symmetries. The relevant condensed matter systems include high-temperature superconductors (cuprates), heavy fermion systems, transition metal oxides, twisted bilayer graphene which exhibit such interaction dominant phases as correlated insulating phases, Wigner crystal, unconventional superconductivity, and non-Fermi liquid behavior [3, 4, 5, 6].
@@
Quasiparticles with the dispersion relation $E(p)\sim p^{2n}$, $n\geq 2$, refer to emergent excitations in quantum many-body systems whose energy-momentum dispersion is soft, meaning that their energy increases slowly with momentum near zero momentum. Such systems naturally realize the regime of dominance of potential energy over the kinetic energy and for large $n$ describe and model weakly dispersing flat bands. Flat bands are especially prone to strong correlation effects as electrons’ kinetic energy is quenched and interactions dominate. Theoretical studies explore how even weakly dispersing flat bands can lead to enhanced susceptibilities, magnetism, superconductivity, or localization [7].
The problem of localization is related to the study of disorder and impurity induced bound states. The latter problem provides us with motivation for the investigation of bound states in systems with quartic dispersion performed in this paper. Mathematically, finding eigenvalues and eigenfunctions of bound states for quasiparticles with soft dispersion implies solving higher order Schrödinger-like equation in the presence of an interaction potential which is not an easy task.
The semiclassical Wentzel-Kramers-Brillouin (WKB) approach is an efficient method for determining approximate solutions to linear differential equations with spatially varying coefficients [8, 9, 10]. In the last decades, the WKB method for the one-dimensional second-order Schrödinger equation has showed a significant progress. One should mention such achievements as the method of the Maslov canonical operator [11], the summation of asymptotic series in the Planck constant $\hbar$ (the exact (complex) WKB method [12, 13]), the finding of ”instanton-like” non-perturbative in $\hbar$ corrections to the Bohr-Sommerfeld quantization condition [14, 15]. The method has been extended to higher spatial dimensions and multi-component Hamiltonians [16] and used to study such physical systems as graphene (see, e.g., Refs.[17, 18, 19]).
Two of us applied recently the WKB method to determine bound state energies for quasiparticles with quartic dispersion [20]. It should be noted that for fourth-order ordinary differential equations with generic potential $V(x)$ only local WKB solutions in low orders in the Planck constant $\hbar$ are available in the literature (see, e.g., [21, 22, 23]), but their global properties, related to quantization, only begin to be investigated [24].
@@
Since the WKB method is not applicable in vicinity of turning points, to determine energies of bound states requires matching semiclassical wave functions in the classically allowed and forbidden regions and finding the connection formulas. The latter could be derived by using solutions to the equation with linearised potential in vicinity of turning points. For the quartic dispersion, such solutions are given by the higher-order Airy functions [25, 26]. Studying its asymptotics on the real axis as $x\to\pm\infty$ by using the method of steepest descents [27, 28] and determining the leading asymptotics as well as the exponentially suppressed terms (hyperasymptotics) necessary for matching the WKB solutions in vicinity of the turning points, the corresponding Bohr-Sommerfeld quantization condition was obtained in [20]. It contains a non-perturbative in $\hbar$ term which is relevant for low-energy states.
Here we extend the study performed in [20] aiming to account for higher order WKB corrections in powers of $\hbar$ that provides an important information on energy levels. In addition, since the WKB expansion is the asymptotic one and the series in $\hbar$ does not capture non-analytic in $\hbar$ contributions, we include non-perturbative contributions due to hyperasymptotics in our analysis. Since, to the best of our knowledge, there are no exact solutions to the double quartic problem (with the Hamiltonian $H=a^{4}\hat{p}^{4}+b^{4}x^{4}$) and to check the validity of our results, we determine the bound state energies numerically by applying the variational approach utilizing the universal Gaussian basis. However, since it is always instructive to have an exact solvable example, we find bound state energies and their wave functions for the exactly solvable problem of a square potential well. One of our main findings is that in the case of quartic dispersion, although the wave functions of bound states in the classically forbidden region exponentially decay, they have nodes in all considered examples. This invalidates the oscillation theorem characteristic for the one-dimensional second-order Schrödinger equation. Still it is observed that the oscillation theorem holds in the classically allowed region.
@@
The paper is organized as follows. The WKB method is formulated in Sec. II. The Wentzel complex method is considered in Sec. III. In Sec.IV, numerical analysis for the bound state energies and wave functions is performed. Bound states for the square well potential are found in Sec.V. The results are summarized in Sec.VI.
II WKB solutions for quartic energy-momentum dispersion
@@
$H=a^{4}p^{4}+V(x),$ (1)
where $p=-i\hbar\partial_{x}$ is the momentum operator, $a^{4}$ is parameter whose dimension is $v^{4}/W^{3}$, where $v$ is velocity and $W$ is energy (cf. with the energy dispersion in ABC-stacked tetra-layer graphene where $E(p)=(v_{F}p)^{4}/\gamma^{3}_{1}$).
To solve the stationary Schrödinger-like equation $H\psi=E\psi$ we use the WKB ansatz for the wave function
$\psi(x)=\exp[iS(x)/\hbar]$ (2)
and obtain the following equation:
$a^{4}\left[(S^{\prime})^{4}-6i\hbar S^{\prime\prime}(S^{\prime})^{2}-\hbar^{2}(4S^{\prime}S^{\prime\prime\prime}+3(S^{\prime\prime})^{2})-i\hbar^{3}S^{IV}\right]+V(x)=E,$ (3)
where primes denote derivatives with respect to $x$. Expanding $S$ in powers of $\hbar$, $S=\sum_{n=0}^{\infty}\hbar^{n}S_{n}$, one can get recursive relations for $S_{n}$. For example, in the zero order, we find the classical momentum
@@
which gives $S_{0}(x)=\int^{x}dx(E-V(x))^{1/4}/a$. Notice that terms in the expansion $S^{\prime}=\sum_{n=0}^{\infty}\hbar^{n}S^{\prime}_{n}$ with odd $n\geq 3$ are total derivatives like in the case of the Schrödinger equation and vanish after integrating over the closed contour in the complex plane encircling turning points [24].
For the first and second order corrections, we have the following equations:
$4(S^{\prime}_{0})^{3}S^{\prime}_{1}-6iS^{\prime\prime}_{0}(S^{\prime}_{0})^{2}=0,$ (5)
$6(S_{0}^{\prime})^{2}(S_{1}^{\prime})^{2}+4(S_{0}^{\prime})^{3}S_{2}^{\prime}-12iS_{0}^{\prime}S_{1}^{\prime}S_{0}^{\prime\prime}-3(S_{0}^{\prime\prime})^{2}-6i(S_{0}^{\prime})^{2}S_{1}^{\prime\prime}-4S_{0}^{\prime}S_{0}^{\prime\prime\prime}=0,$ (6)
which give
$S^{\prime}_{1}=\frac{3iS^{\prime\prime}_{0}}{2S^{\prime}_{0}}=\frac{3i}{2}\left(\ln S^{\prime}_{0}\right)^{\prime}=\frac{3i}{2}\left(\ln p(x)\right)^{\prime},$ (7)
$S^{\prime}_{2}=\frac{5}{4}\left(\frac{3(p^{\prime})^{2}}{2p^{3}}-\frac{p^{\prime\prime}}{p^{2}}\right)=-\frac{5}{4}\left(\frac{(p^{\prime})^{2}}{2p^{3}}+\left(\frac{p^{\prime}}{p^{2}}\right)^{\prime}\right),$ (8)
where we used $S^{\prime}_{0}(x)=p(x)$. Integrating the equation for $S_{1}^{\prime}$, we obtain $S_{1}=i\ln p^{3/2}$.
The fourth order correction $S_{4}^{\prime}$ is given in Ref.[24] (up to total derivatives) and has the following form:
$S_{4}^{\prime}=\frac{35(p^{\prime\prime})^{2}}{256\,p^{5}}-\frac{21p^{IV}}{1024\,p^{4}}.$ (9)
The higher order terms with $n\geq 6$ can be also obtained recursively and expressed through $p(x)$ and its derivatives (for terms $S^{\prime}_{6},S^{\prime}_{8}$, see Ref.[24]). However, in this paper, we restrict ourselves to the corrections up to the 4th order. In the next section, we use the obtained results and apply the Wentzel complex method to determine the bound state energies for the system with the quartic dispersion and the quartic potential $V(x)=b^{4}x^{4}$.
III Quantization condition for double quartic problem
The Wentzel complex method uses the single-valuedness of a wave function $\psi$ in the region containing turning points. The argument principle then leads to the condition
@@ III Quantization condition for double quartic problem
where $n$ is an integer number and the contour $\gamma$ encloses the classical turning points anticlockwise in the complex $z$-plane. For quartic dispersion and in the fourth order in $\hbar$, we have
@@
Finally, we obtain the following equation for the WKB energy levels of the double quartic problem:
$\frac{\Gamma^{2}(1/4)\sqrt{\varepsilon}}{4\sqrt{\pi}}-\frac{3\Gamma^{2}(3/4)}{4\sqrt{\pi}}\frac{1}{\sqrt{\varepsilon}}-\frac{3\Gamma^{2}(1/4)}{128\sqrt{\pi}\varepsilon^{3/2}}=\pi\left(n+\frac{1}{2}\right),\qquad n=0,1,\dots\,.$ (17)
Adding non-perturbative contribution due to hyperasymptotics derived in Ref.[20], we obtain the quantization condition
$\frac{\Gamma^{2}(1/4)\sqrt{\varepsilon}}{4\sqrt{\pi}}-\frac{3\Gamma^{2}(3/4)}{4\sqrt{\pi}}\frac{1}{\sqrt{\varepsilon}}-\frac{3\Gamma^{2}(1/4)}{128\sqrt{\pi}\varepsilon^{3/2}}-2(-1)^{n}\arctan\left[\frac{1}{2}\exp\left(-\frac{\Gamma^{2}(1/4)\sqrt{\varepsilon}}{4\sqrt{\pi}}\right)\right]=\pi\left(n+\frac{1}{2}\right),\qquad n=0,1,\dots\,.$ (18)
@@
Table 1: Numerical values of the bound state energy $\varepsilon$ for the one-dimensional system with quartic dispersion and quartic potential for $n=0,1,\dots,6$ taking into account the WKB corrections of the second $\varepsilon_{2ndcorr}$ and fourth $\varepsilon_{4thcorr}$ order, as well as the non-perturbative correction due to hyperasymptotics.
Numerical computation of the WKB energies for the double quartic system defined by Eqs.(17) and (18) are presented in Table 1 for the states with n=0,1,,6n=0,1,...,6. The second and fourth rows of the table show the WKB energies with the second and fourth order WKB corrections taken into account, respectively. To quantify the role of the non-perturbative contribution, we include in the third and fifth rows the non-perturbative in $\hbar$ contribution given by the last term on the left-hand side of Eq.(18). The last row presents variational estimates which are derived in the next section.
@@
IV Numerical results for the lowest energy levels
To check the validity and accuracy of the results obtained in the previous section, we determine bound state energies for the harmonic $V\left(x\right)=x^{2}$ and quartic $V\left(x\right)=x^{4}$ potentials using the variational method with the universal Gaussian basis [30].
@@ IV Numerical results for the lowest energy levels
Let us consider now the wave functions of the bound states. Using Eq.(19) and the determined coefficients $C_{k}$ and parameters $a_{k}$, $x_{k}$, we depict the bound state wave functions of the ground and first excited states for the Hamiltonian $H=p^{4}+x^{4}$ in the left and middle panels of Fig.2 (red solid lines). For comparison, we show in the same figure the wave functions for the harmonic potential $V\left(x\right)=x^{2}$ (blue dashed lines). Energies of the ground and first excited states for the Hamiltonian $H=p^{4}+x^{4}$ are $E=1.3967$ and $E=7.1315$, respectively. For the Hamiltonian with quadratic potential $H=p^{4}+x^{2}$, the corresponding energies are lower and equal $E=1.0603$ and $E=3.7996$. Vertical red solid and blue dashed lines in the left and middle panels show boundaries of the classically allowed region for the ground and first excited states, respectively. While for the ground state of the Hamiltonian $H=p^{4}+x^{4}$ the classically allowed region is defined by 1.0871<x<1.0871-1.0871<x<1.0871 for the ground state, it is given by 1.6341<x<1.6341-1.6341<x<1.6341 for the first excited state. As to the classically allowed region for the Hamiltonian $H=p^{4}+x^{2}$, it is 1.0297<x<1.0297-1.0297<x<1.0297 and 1.94927<x<1.94927-1.94927<x<1.94927 for the ground and first excited states, respectively. The wave functions are normalized to the unity, i.e., $\int\left|\psi\left(x\right)\right|^{2}dx=1$. Oscillations of the wave functions of the ground and first excited states for the potential $V(x)=x^{4}$ in the classically forbidden region are shown in insets of all panels of Fig.2. It is instructive to compare the wave functions of the ground and first excited states of the double quartic problem with the familiar case of the quadratic energy-momentum dispersion. The wave functions of the later problem are plotted in the right panel of Fig.2.
@@
Figure 2: Left panel: The ground state wave function for the $H=\hat{p}^{4}+x^{4}$ (red solid line) and $H=\hat{p}^{4}+x^{2}$ Hamiltonians (blue dashed line). Middle panel: The wave functions of the first excited state for the $H=\hat{p}^{4}+x^{4}$ (red solid line) and $H=\hat{p}^{4}+x^{2}$ Hamiltonians (blue dashed line). Vertical red solid and blue dashed lines in the left and middle panels show boundaries of the classically allowed region for the ground and first excited states, respectively. Right panel: The wave function of the ground state (orange solid line) and the first excited state (brown dashed line) for the quartic oscillator with the Hamiltonian $H=\hat{p}^{2}+x^{4}$.
@@
V Finite square well potential

### v2

# Nodal structure of bound-state wave functions for systems with quartic dispersion

@@ Abstract
The nodal structure of bound-state wave functions for one-dimensional quantum systems with quartic energy–momentum dispersion, $E(p)\sim p^{4}$, and polynomial potentials is analysed by using the semiclassical approximation and variational approach. For energies of bound states, we derive the quantization condition, obtained by using the complex Wentzel method, where we take into account perturbative (up to the fourth order) and nonperturbative in the Planck constant corrections. The bound-state energies and wave functions for the harmonic and quartic potentials are compared with those found by applying the variational approach utilizing the universal Gaussian basis. It is shown that the classical oscillation theorem, valid for systems with quadratic energy-momentum dispersion, breaks down in the classically forbidden region where wave functions also have nodes, while it still remains valid in the classically allowed region. These results are confirmed in addition via the solutions of the exactly solvable problem of the fourth-order Schrödinger equation with a square well potential.
@@ I Introduction
Quantum mechanical systems in which potential energy dominates over the kinetic energy attract a significant interest because they often exhibit strongly correlated behavior and give rise to novel quantum phases of matter which are not well described by traditional perturbative or mean-field approaches [1, 2]. These systems are central to many open questions in modern physics including entanglement, the simulation of strongly interacting gauge theories in condensed matter analogs, and the realization of quantum criticality and emergent symmetries. The relevant condensed matter systems include high-temperature superconductors (cuprates), heavy fermion systems, transition metal oxides, twisted bilayer graphene which exhibit such interaction dominant phases as correlated insulating phases, Wigner crystal, unconventional superconductivity, and non-Fermi liquid behavior [3, 4, 5, 6].
@@
Quasiparticles with the dispersion relation $E(p)\sim p^{2n}$, $n\geq 2$, refer to emergent excitations in quantum many-body systems whose energy-momentum dispersion is soft, meaning that their energy increases slowly with momentum near zero momentum. Such systems naturally realize the regime of dominance of potential energy over the kinetic energy and for large $n$ describe weakly dispersing bands. Theoretical studies explore how even weakly dispersing bands can lead to enhanced susceptibilities, magnetism, superconductivity, or localization [7].
The problem of localization is related to the study of disorder and impurity induced bound states thus providing a motivation for investigating bound states in systems with quartic dispersion. An important characteristics of bound states is the number of nodes of their wave functions. The nodal structure of wave functions of bound states for systems with quartic dispersion is the main focus of the present paper.
@@
Mathematically, finding eigenvalues and eigenfunctions of bound states for quasiparticles with soft dispersion implies solving higher order Schrödinger-like equation in the presence of an interaction potential which is not an easy task. Recently, two of us applied the semiclassical Wentzel-Kramers-Brillouin (WKB) approach to determine bound state energies for quasiparticles with quartic dispersion and polynomial potentials [8]. It should be noted that for fourth-order ordinary differential equations with generic potential $V(x)$ only local WKB solutions in low orders in the Planck constant $\hbar$ are available in the literature (see, e.g., [9, 10, 11]), but their global properties, related to quantization, only begin to be investigated [12]. Since the WKB method is not applicable in vicinity of turning points, to match semiclassical wave functions in the classically allowed and forbidden regions one needs to obtain the connection formulas. The latter are defined by solutions to the equation with linearised potential in vicinity of turning points which are given by the fourth-order Airy functions [13, 14] in the case of quartic dispersion. Determining the leading asymptotics for these functions and their exponentially suppressed terms (hyperasymptotics) necessary for matching the WKB solutions in vicinity of the turning points, the corresponding Bohr-Sommerfeld quantization condition was obtained in [8] which contains a non-perturbative in $\hbar$ term related to hyperasymptotics.
Here we use the results obtained in [8] for the harmonic and quartic potentials to analyse the energies of bound states accounting also for higher order (up to $\hbar^{4}$) WKB corrections. The emphasis is placed on the potential of an anharmonic oscillator $V(x)\sim x^{4}$, which has been extensively studied in quantum mechanics for many years (see, book [15] and the recent article [16] with numerous references therein). Since, to the best of our knowledge, there are no exact solutions to the double quartic problem (with the Hamiltonian $H=a^{4}\hat{p}^{4}+b^{4}x^{4}$) and to check the WKB results, we determine the bound state energies numerically by applying the variational approach utilizing the universal Gaussian basis. In addition, it is always instructive to have an exact solvable example, therefore, we find bound state energies and their wave functions for the exactly solvable problem of a square potential well.
One of our main findings is that in the case of quartic dispersion, although the wave functions of bound states in the classically forbidden region exponentially decay, they have nodes in that region in all considered examples. This invalidates the oscillation theorem characteristic for the one-dimensional second-order Schrödinger equation. The latter states that the $n$th eigenenergy bound-state wave function of the 1D Schrödinger equation has exactly $n$ nodes within classically allowed region and no nodes in classically forbidden region. Still it is observed that the oscillation theorem holds in the classically allowed region for the fourth-order Schrödinger equation.
@@
The paper is organized as follows. The WKB approximation and the Wentzel complex method are considered in Sec. II for the double quartic problem. In Sec.III, bound state energies and wave functions are analysed using the variational method. Bound states for the square well potential are found in Sec.IV. The results are summarized in Sec.V.
II Quantization condition for double quartic problem
@@
$H=a^{4}\hat{p}^{4}+V(x),$ (1)
where $\hat{p}=-i\hbar\partial_{x}$ is the momentum operator, $a^{4}$ is parameter whose dimension is $v^{4}/W^{3}$, where $v$ is velocity and $W$ is energy (compare with the energy dispersion in tetra-layer graphene where $E(p)=(v_{F}p)^{4}/\gamma^{3}_{1}$ [17, 18, 19]). We assume the potential $V(x)$ to be a continuous function which tends to infinity for $|x|\to\infty$ and has a single global minimum which ensures the existence of two real turning points.
To solve the stationary Schrödinger-like equation $H\psi=E\psi$ we use the WKB ansatz for the wave function $\psi(x)=\exp[iS(x)/\hbar]$ and obtain the following equation:
$a^{4}\left[(S^{\prime})^{4}-6i\hbar S^{\prime\prime}(S^{\prime})^{2}-\hbar^{2}(4S^{\prime}S^{\prime\prime\prime}+3(S^{\prime\prime})^{2})+i\hbar^{3}S^{IV}\right]+V(x)=E,$ (2)
where primes denote derivatives with respect to $x$. Expanding $S$ in powers of $\hbar$, $S=\sum_{n=0}^{\infty}\hbar^{n}S_{n}$, one can get recursive relations for derivatives of $S_{n}$ (for some low order recursive relations see Ref.[8]). For example, in the zero order, we find the classical momentum
@@
which gives $S_{0}(x)=\int^{x}dx(E-V(x))^{1/4}/a$. From recursive relations we find for $S^{\prime}_{n}$ the following expressions up to $n=4$:
$S^{\prime}_{1}=\frac{3i}{2}\left(\ln p(x)\right)^{\prime},\quad S^{\prime}_{2}=\frac{5}{4}\left(\frac{3(p^{\prime})^{2}}{2p^{3}}-\frac{p^{\prime\prime}}{p^{2}}\right)=-\frac{5}{4}\left(\frac{(p^{\prime})^{2}}{2p^{3}}+\left(\frac{p^{\prime}}{p^{2}}\right)^{\prime}\right),$ (4)
$S^{\prime}_{3}=\frac{5i}{8}\left(\frac{3(p^{\prime})^{2}}{2p^{4}}-\frac{p^{\prime\prime}}{p^{3}}\right)^{\prime},\qquad S^{\prime}_{4}=\frac{35(p^{\prime\prime})^{2}}{256\,p^{5}}-\frac{21p^{IV}}{1024\,p^{4}}+\left(\frac{75(p^{\prime})^{3}}{256p^{6}}-\frac{75p^{\prime}p^{\prime\prime}}{256p^{5}}+\frac{85p^{\prime\prime\prime}}{1024p^{4}}\right)^{\prime}.$ (5)
Integrating the equation for $S_{1}^{\prime}$, we obtain $S_{1}=i\ln p^{3/2}$. Notice that terms in the expansion $S^{\prime}=\sum_{n=0}^{\infty}\hbar^{n}S^{\prime}_{n}$ with odd $n\geq 3$ are total derivatives like in the case of the Schrödinger equation and vanish after integrating over the closed contour in the complex plane encircling turning points [12]. The total derivatives terms in $S^{\prime}_{n}$ with even $n\geq 2$ do not contribute either.
The higher order terms with $n\geq 6$ can be also obtained recursively and expressed through $p(x)$ and its derivatives (for terms $S^{\prime}_{6},S^{\prime}_{8}$, see Ref.[12]). However, the algebraic complexity of expressions increases rapidly with increasing $n$ and, at a certain stage, requires the use of a computer. In this paper, we restrict ourselves to the corrections up to the 4th order. We now use the obtained results and apply the Wentzel complex method to determine the bound state energies for the system with the quartic dispersion and the quartic potential $V(x)=b^{4}x^{4}$.
The Wentzel complex method uses the single-valuedness of a wave function $\psi(z)$ in the region of complex plane $z$ containing turning points. Indeed, on the real axis the bound state wave function is a real function and must decrease in the classically forbidden region while oscillating in the classically allowed region producing nodes. To count for nodes due to oscillations the argument principle should be used which then leads to the condition
@@
where $n$ is an integer number and the contour $\gamma$ encloses the classical turning points anticlockwise in the complex $z$-plane. Since the semiclassical momentum is a fourth-root expression resulting in a four-sheeted Riemann surface of complex variable $z$, the contour should be chosen so that the integrand $S^{\prime}(z)$ return to its original value after going along the contour. Encompassing a turning point, the momentum acquires the phase factor $e^{i\pi/2}$, therefore, one needs to go around each turning point twice, as shown in Fig.1, to return to the original value of the momentum. Since the integrand term $S^{\prime}_{n}(z)$, after calculating the derivatives, contains only integer powers of $z$ and $p(z)$, it is a single-valued function on the considered contour. In addition, the contour should be drawn close to the real axis avoiding possible zeros in the complex domain, as well as zeros in the classically forbidden domain, because we need to account for nodes of a bound state wavefunction on the real axis between turning points. Nodes in the classically forbidden region, as will be shown below, appear in eigenfunctions of the fourth-order differential operator for which, in contrast to the Schrödinger operator, the oscillation theorem does not hold in its standard formulation.
For quartic dispersion and in the fourth order in $\hbar$, we have
@@
Finally, we obtain the following quantization condition which defines the WKB energy levels of the double quartic problem:
$\frac{\Gamma^{2}(1/4)\sqrt{\varepsilon}}{4\sqrt{\pi}}-\frac{3\Gamma^{2}(3/4)}{4\sqrt{\pi}}\frac{1}{\sqrt{\varepsilon}}-\frac{3\Gamma^{2}(1/4)}{128\sqrt{\pi}\varepsilon^{3/2}}-2(-1)^{n}\arctan\left[\frac{1}{2}\exp\left(-\frac{\Gamma^{2}(1/4)\sqrt{\varepsilon}}{4\sqrt{\pi}}\right)\right]=\pi\left(n+\frac{1}{2}\right),\qquad n=0,1,\dots\,$, (13)
where the last term on the left-hand side of the above equation accounts for the non-perturbative contribution due to hyperasymptotics derived in Ref.[8].
@@
Table 1: Numerical values of the bound state energy $\varepsilon$ for the one-dimensional system with quartic dispersion and quartic potential for $n=0,1,\dots,6$ taking into account the WKB corrections of the second and fourth order $\varepsilon_{4thcorr}$, as well as the non-perturbative correction due to hyperasymptotics. Bound-state energies obtained using the variational approach in Sec.III are shown in the last row.
Numerical results of computation of the WKB energies for the double quartic system defined by Eq.(13) are presented in Table 1 for the states with n=0,1,,6n=0,1,...,6. The second row shows the WKB energies with the second and fourth order WKB corrections taken into account. To quantify the role of the non-perturbative contribution, we include in the third row the non-perturbative in $\hbar$ contribution given by the last term on the left-hand side of Eq.(13). The last row presents variational estimates which are derived in the next section.
@@
III Variational method and nodal structure of bound-state wave functions
To analyse the properties of bound-state wave functions and check the validity and accuracy of the results obtained in the previous section, we apply here the variational method with the universal Gaussian basis [21] for systems with quartic energy-momentum dispersion and the harmonic $V\left(x\right)\sim x^{2}$ and quartic $V\left(x\right)\sim x^{4}$ potentials.
@@ III Variational method and nodal structure of bound-state wave functions
Let us consider now the wave functions of the bound states. Using Eq.(14) and the determined coefficients $C_{k}$ and parameters $a_{k}$, $x_{k}$, we depict the bound state wave functions of the ground and first excited states for the Hamiltonian $H=\hat{p}^{4}+x^{4}$ in the left and middle panels of Fig.2 (red solid lines). For comparison, we show in the same figure the wave functions for the harmonic potential $V\left(x\right)=x^{2}$ (blue dashed lines). Energies of the ground and first excited states for the Hamiltonian $H=\hat{p}^{4}+x^{4}$ are $E=1.3967$ and $E=7.1315$, respectively. For the Hamiltonian with quadratic potential $H=\hat{p}^{4}+x^{2}$, the corresponding energies are lower and equal $E=1.0603$ and $E=3.7996$. Vertical red solid and blue dashed lines in the left and middle panels show boundaries of the classically allowed region for the ground and first excited states, respectively. While for the ground state of the Hamiltonian $H=\hat{p}^{4}+x^{4}$ the classically allowed region is defined by 1.0871<x<1.0871-1.0871<x<1.0871 for the ground state, it is given by 1.6341<x<1.6341-1.6341<x<1.6341 for the first excited state. As to the classically allowed region for the Hamiltonian $H=\hat{p}^{4}+x^{2}$, it is 1.0297<x<1.0297-1.0297<x<1.0297 and 1.94927<x<1.94927-1.94927<x<1.94927 for the ground and first excited states, respectively. The wave functions are normalized to the unity, i.e., $\int\left|\psi\left(x\right)\right|^{2}dx=1$. Oscillations of the wave functions of the ground and first excited states for the potential $V(x)=x^{4}$ in the classically forbidden region are shown in insets of all panels of Fig.2. It is instructive to compare the wave functions of the ground and first excited states of the double quartic problem with the familiar case of the quadratic energy-momentum dispersion. The wave functions of the later problem are plotted in the right panel of Fig.2.
@@
Figure 2: Left panel: The ground state wave function for the $H=\hat{p}^{4}+x^{4}$ (red solid line) and $H=\hat{p}^{4}+x^{2}$ Hamiltonians (blue dashed line). Middle panel: The wave functions of the first excited state for the $H=\hat{p}^{4}+x^{4}$ (red solid line) and $H=\hat{p}^{4}+x^{2}$ Hamiltonians (blue dashed line). Vertical red solid and blue dashed lines in the left and middle panels show boundaries of the classically allowed region for the ground and first excited states. Right panel: The wave function of the ground state (orange solid line) and the first excited state (brown dashed line) for the quartic oscillator with the Hamiltonian $H=\hat{p}^{2}+x^{4}$.
@@
IV Spectrum and wave function nodes for square well potential